A monopoly is considering selling several units of a homogeneous product as a single package. A typical consumer’s demand for the product is Qd = 40 - 0.5P, and the marginal cost of production is $50.

a. Determine the optimal number of units to put in a package.

_____ units

b. How much should the firm charge for this package?

$_____
Answer #1

a) QD=40-0.5P

So, P=80-2Q

The optimal number of units is produced where P=MC

80-2Q=50

80-50=2Q

Q=30/2 = 15

b) Price charged = CS+TR

CS=0.5*15*(80-50) = 225

TR=15*50 = 750

Price charged = 750+225 = 975